# Integration of Cobalt Ferromagnetic Control Gates for Electrical and Magnetic Manipulation of Semiconductor Quantum Dots

### v1

@@ 3.2 Nanogates Magnetization Pattern
where r is the position vector, the -component of the magnetic vector potential, the elementary charge, and the reduced Planck’s constant. The electron holography measurement allows mapping the in-plane components of magnetic induction (i.e., ) integrated along the thickness of the sample in the Coulomb gauge (, where is the nabla operator and the cross product, see SI for details). Mapping of the magnetic phase and the magnetic induction (), integrated over the out-of-plane direction, are reported in Figure 6. As expected, the measured remanent magnetization does not align with the magnetization axis, and an intercoupling between two magnetic domains is observed in the larger Co pad. To investigate the role of polycrystallinity, micromagnetic simulations with a 20 nm sized crystallite random pattern were performed. The simulation is run on a mesh of polyhedra defined using the method of Voronoi tesselation and assigns random crystal directions to each polyhedron, mimicking the non-textured polycrystalline structure shown by XRD and TEM. Figure 6b shows one simulation outcome with a magnetization pattern that leads to a closely resembling the holography experiment, at least for the middle and rightmost magnetic pad. The vortex-like structures visible in the middle and rightmost magnetic pads minimize the stray field, but the overall complicated magnetic distribution suggests that the crystallite texture and fabrication-induced defects play a major role in determining the local magnetization orientation. Indeed, simulations which differ only by a different crystallite pattern do not lead to similar results (see SI). This suggests investigating a lower bound for the external magnetic field , knowing that the requirements for an effective control of spin states through EDSR (large driving gradients, small dephasing gradients and single qubit addressability [11]) demand uniform magnetization along the nanowire axis [12]. The minimum value of to enforce uniform magnetization is investigated by micromagnetic simulations, where the nanogates shape wrapping the Si nanowire is taken into account as well as the polycrystalline structure of Co (SI). Figure 7a shows that T ensures an almost homogeneous magnetization within the nanogates, setting a lower boundary for , whereas at lower magnetic field ( T) the magnetic shape anisotropy and magnetocrystalline anisotropy dominate the magnetization pattern orientation. This leads to a non-uniform stray field in the qubit plane, as shown in Figure 7b. Thus, we conclude that the fabricated Co nanogates are suitable for EDSR drive of spin qubits at large Larmor frequencies ( T), where homogeneous magnetization is ensured along the hard magnetic axis of the nanostructures.
@@ 4 Conclusion
In this study we analyzed the integration of Co nanogates for electrical and magnetic manipulation of quantum dots in Si and validated its compatibility with FEOL CMOS processes. We examined the morphology of e-beam evaporated thin Co films using ARXPS and GIXRD, observing interlayer diffusion of the adhesion layer (Cr) into Co after thermal treatment in forming gas at 300 °C, along with limited surface oxidation in samples that were carefully cooled in forming gas. XPS spectra confirmed the formation of CoO/Co2O3/Co3O4 and Co(OH2) at the surface of the sample. Under our process conditions, the partial oxidation was limited to 4 nm and did not impact the magnetic properties of the underlying Co, as measured by VSM. Electron diffraction patterns indicate the formation of an fcc phase for annealing temperatures above 300 °C, and no evolution of the grain size was measured within the temperature range used in this work. We compared the work function and charge traps of Ti-Pd and Cr-Co gate stacks by performing high-low frequency characterization of MOS capacitors, noting a 15 increase in effective work function of the ferromagnetic gate-stack and a reduction of charge traps to cm-3 when thermally annealed in forming gas at 300 °C for 10 minutes. No relevant difference of interface traps was measured between Ti/Pd and Cr/Co gate stacks. To validate the process, we fabricated an FDSOI nanowire device featuring Pd and Co control gates for the electrostatic control of quantum dots. Electrical characterizations at room temperature and 10 mK confirmed accurate field-effect control of the charge density in SOI, and stable quantum confinement was observed at cryogenic temperature for Co gates biased below threshold. Electron holography measurements coupled with micromagnetic simulations suggest that local magnetic texture inside the ferromagnetic gates is mostly controlled by the crystallite structure and fabrication-induced defects, both dependent on thermal treatments and material morphology. Micromagnetic simulations mimicking the non-textured polycrystalline structure of Co measured by XRD and TEM have been used to determine a lower bound for the external magnetic field to ensure homogeneous magnetization along the hard axis of the ferromagnetic gates. Simulation results show that for magnetic fields lower than 1 T, the magnetic shape and magnetocrystalline anisotropy lead to a non-uniform stray field in the plane of the quantum dots, thus establishing a minimum operation condition for the EDSR drive. These findings demonstrate the feasibility of FEOL/MOL-compatible integration of Co in silicon-based quantum devices, with emphasis on ferromagnetic control gates suitable for both electrostatic control and EDSR driving of electron spin qubits.
@@ 5 Experimental Section
Samples Fabrication: The SOI nanowires and all layer of gates were patterned by electron beam lithography using a Raith EPBG5000+ system with HSQ and PMMA 950K resists. The substrates were provided by Soitec and thinned down to 18 nm by several cycles of dry oxidation in a Centrotherm furnace and wet etching in buffered HF. All metals were deposited through e-beam evaporation and patterned by lift-off in acetone. Low power oxygen plasma was used to remove any residues of polymers. The aluminum oxide was deposited by ALD at 300 °C using Trimethylalluminum and H2O precursors in a Beneq TFS200 system. The TEM lamellae were prepared by Ga FIB and stored under vacuum.
Material Characterization: GIXRD analysis was performed on a Malvern Panalytical Empyrean thin films diffractometer, equipped with a monochromated Cu K source and a PIXcel3D detector. XPS and ARXPS analyses were performed using a Kratos Axis Supra equipped with a monochromated Al K source. SEM EDX imaging was conducted using a Zeiss Merlin SEM equipped with a X-Max 50 mm2 detector from Oxford Instruments. TEM imaging, EDX and SAED measurements were performed using a FEI Tecnai Osiris microscope, operated at 200 kV and equipped with a Gatan Orius camera. The VSM measurements were performed with a Lake Shore Cryotronics 7300 Series Vibrating Sample Magnetometer. Hysteresis loops were measured by applying the external field parallel to the surface of the thin film (i.e. within the easy plane set by shape anisotropy). Electron holography measurements were performed on a FEI Titan Themis microscope operated at 300 kV, equipped with an electron biprism and a Gatan K3 camera; -phase shifting off-axis holography method was used [60] and electrostatic contribution was removed to access the magnetic contribution.
@@
Supporting information for XRD and XPS analysis, electron holography measurements, and micromagnetic simulations is available upon request from the authors.
@@ Acknowledgements
The authors acknowledge the Center of Micro-NanoTechnology (CMi) and the Interdisciplinary Center for Electron Microscopy (CIME) of EPFL for their support during the development of the fabrication process and characterization of the materials. A special gratitude goes to Dr. Rolf Allenspach (IBM Research-Zurich) for his valuable insights on the micromagnetic simulations, as well as to Dr. Floris Braakman, Liza Žaper, and prof. Martino Poggio from the University of Basel for their insightful discussions regarding the presented architecture. This work was supported as a part of NCCR SPIN, a National Centre of Competence in Research, funded by the Swiss National Science Foundation (grant number 225153).

### v2

@@ 3.2 Nanogates Magnetization Pattern
where r is the position vector, the -component of the magnetic vector potential, the elementary charge, and the reduced Planck’s constant. The electron holography measurement allows mapping the in-plane components of magnetic induction (i.e., ) integrated along the thickness of the sample in the Coulomb gauge (, where is the nabla operator and the cross product, see SI for details). Mapping of the magnetic phase and the magnetic induction (), integrated over the out-of-plane direction, are reported in Figure 6. As expected, the measured remanent magnetization does not align with the magnetization axis, and an intercoupling between two magnetic domains is observed in the larger Co pad. To investigate the role of polycrystallinity, micromagnetic simulations with a 20 nm sized crystallite random pattern were performed. The simulation is run on a mesh of polyhedra defined using the method of Voronoi tesselation and assigns random crystal directions to each polyhedron, mimicking the non-textured polycrystalline structure shown by XRD and TEM. Figure 6b shows one simulation outcome with a magnetization pattern that leads to a closely resembling the holography experiment, at least for the middle and rightmost magnetic pad. The vortex-like structures visible in the middle and rightmost magnetic pads minimize the stray field, but the overall complicated magnetic distribution suggests that the crystallite texture and fabrication-induced defects play a major role in determining the local magnetization orientation. Indeed, simulations which differ only by a different crystallite pattern do not lead to similar results (see SI). This suggests investigating a lower bound for the external magnetic field , knowing that the requirements for an effective control of spin states through EDSR (large driving gradients, small dephasing gradients and single qubit addressability [11]) demand uniform magnetization along the nanowire axis [12].
3.2.1 EDSR Driving by Co Nanogates
The minimum value of to enforce uniform magnetization is investigated by micromagnetic simulations, where the shape of the nanogates wrapping the Si nanowire is taken into account, as well as the polycrystalline structure of Co (see SI).
Figure 7a shows that T ensures an almost homogeneous magnetization within the nanogates, whereas at T the magnetic shape anisotropy and magnetocrystalline anisotropy dominate the magnetization pattern orientation. This leads to a non-uniform stray field in the qubit plane for T, as shown in Figure 7b, with the result of an impredictable addressability (i.e., difference in Larmor frequency) of the two adjacent qubits. Assuming sufficiently high fields, a strong magnetic field gradient along the direction normal to the quantization axis () can be created inside the nanowire due to the reduced distance between the ferromagnets and the semiconductor. For qubit operation, the driving gradient is , while dephasing is mainly caused by , since the field symmetry induced by the wrapping gates makes negligible at the qubit plane. We report the simulated gradients generated by the fabricated Co nanogates at low and high field in Figure 7c, considering a magnet-to-dot distance () defined by the gate oxide (with nm) or the gate and interlayer oxide (with nm), depending on whether the Co nanogates are patterned as the first or second metallization layer. At T, the unsaturated magnetization along results in a distorted gradient shape. However, at T, driving gradients exceeding 10 mT/nm are achieved in both configurations, with the decoherence gradient being minimized at the location of the quantum dots. In this experimental condition, the predicted Larmor frequency difference (, assuming a g-factor ) between the two qubits is GHz and GHz for the two gate-stack configurations, respectively. This is beneficial in terms of qubit fidelity, since a fast-driving region coincides with a dephasing protected spot along the main displacement direction, with more than two times larger than the estimated Rabi frequency [12]. Based on these micromagnetic simulation results, we conclude that the fabricated Co nanogates are suitable for EDSR drive of spin qubits at high Zeeman fields only ( T), where homogeneous magnetization is ensured along the hard magnetic axis of the nanostructures.
These considerations, based on numerical simulations, are independent of the dielectric used, assuming a fixed oxide thickness and a consistent distance between the ferromagnets and quantum dots. In the proof-of-concept devices presented in this study, we observed significant charge noise in gate-stack utilizing exclusively Pd gates, as well as in configurations combining Pd and Co control gates. This charge noise, which we attribute to the ALD-Al2O3/Si interface, impeded the precise tuning of double-dot charge configurations required to map a Rabi Chevron pattern via EDSR and Pauli spin blockade. As a result, the demonstration of electron spin qubit operation within the proposed scalable EDSR architecture is deferred to future studies utilizing high-temperature thermal SiO2 in place of ALD-based dielectrics.
@@ 4 Conclusion
In this study we analyzed the integration of Co nanogates for electrical and magnetic manipulation of quantum dots in Si and validated its compatibility with FEOL CMOS processes. We examined the chemistry and morphology of e-beam evaporated thin Co films using ARXPS, GIXRD, and EDX-TEM. Our observations revealed interlayer diffusion of the adhesion layer (Cr) into Co after thermal treatment in forming gas at 300 °C, along with limited surface oxidation in samples that were carefully cooled in nitrogen. XPS spectra confirmed the formation of CoO/Co2O3/Co3O4 and Co(OH2) at the surface of the sample. Under our process conditions, the partial oxidation was limited to 4 nm and did not impact the magnetic properties of the underlying Co, as measured by VSM. Electron diffraction patterns indicate the formation of an fcc phase for annealing temperatures above 300 °C, and limited evolution of the grain size was measured within the temperature range used in this work. We compared the work function and charge traps of Ti-Pd and Cr-Co gate stacks by performing high-low frequency characterization of MOS capacitors, noting a 15 increase in effective work function of the ferromagnetic gate-stack and a reduction of charge traps to cm-3 when thermally annealed in forming gas at 300 °C for 10 minutes. No relevant difference of interface traps was measured between Ti/Pd and Cr/Co gate stacks. To validate the process, we fabricated an FDSOI nanowire device featuring Pd and Co control gates for the electrostatic control of quantum dots. Electrical characterizations at room temperature and 10 mK confirmed accurate field-effect control of the charge density in SOI, and stable quantum confinement was observed at cryogenic temperature for Co gates biased below threshold. Electron holography measurements coupled with micromagnetic simulations suggest that local magnetic texture inside the ferromagnetic gates is mostly controlled by the crystallite structure and fabrication-induced defects, both dependent on thermal treatments and material morphology. Micromagnetic simulations mimicking the non-textured polycrystalline structure of Co measured by XRD and TEM have been used to determine a lower bound for the external magnetic field to ensure homogeneous magnetization along the hard axis of the ferromagnetic gates. Simulation results show that for magnetic fields lower than 1 T, the magnetic shape and magnetocrystalline anisotropy lead to a non-uniform stray field in the plane of the quantum dots which ultimately degrades the control over driving and dephasing gradients, thus establishing a minimum operation condition for the EDSR drive. These findings demonstrate the feasibility of FEOL/MOL-compatible integration of Co in silicon-based quantum devices, with emphasis on ferromagnetic control gates suitable for both electrostatic control and EDSR driving of electron spin qubits.
@@ 5 Experimental Section
Samples Fabrication: The SOI substrates were provided by Soitec and thinned down to 18 nm by several cycles of dry oxidation in a Centrotherm furnace and wet etching in buffered HF. The SOI nanowires and all metal layers were patterned by electron beam lithography using a Raith EPBG5000+ system. A commercial hydrogen silsesquioxane solution (HSQ XR-1541 2) was spin-coated on the SOI sample at 7000 rpm to achieve a uniform thickness of approximately 50 nm. The resist was exposed to a dose of C/cm2 and developed in 25 TMAH at room temperature for 50 s to define a hard mask for the silicon nanowire. The nanowire was then anisotropically etched in an inductively coupled deep reactive ion etching system (Alcatel AMS200) using an SOI-optimized continous etch recipe comprising SF6 and C4F8, with the selectivity to HSQ estimated to be 20:1. The hard mask was then removed by a s dip in HF , given the much higher etch rate compared to the thermal SiO2 forming the buried oxide. A 9 nm Al2O3 gate oxide was deposited by ALD at 300 °C using trimethylalluminum and H2O precursors in a Beneq TFS200 system. Both metallic layers were patterned by lift-off using a single layer of PMMA 950K with an approximate thickness of 90 nm. All metals were deposited via e-beam evaporation: 3 nm for adhesion layers (Ti, Cr) and 33 nm for Pd and Co. The evaporation rate of Co was set to 4 /s at a pressure lower than 1.5 mbar, and the temperature reached during deposition was monitored to remain below 80 °C. Low power oxygen plasma was used to remove any residues of polymers before metals evaporation. A 7 nm ALD-deposited aluminum oxide layer served as the interlayer insulator. Finally, all contacts were opened by maskless laser lithography and wet etching in BHF to access metallic pads for wire-bonding.
The TEM lamellae were prepared by Ga FIB and stored under vacuum.
Material Characterization: GIXRD analysis was performed on a Malvern Panalytical Empyrean thin films diffractometer, equipped with a monochromated Cu K source and a PIXcel3D detector. XPS and ARXPS analyses were performed using a Kratos Axis Supra equipped with a monochromated Al K source. SEM EDX imaging was conducted using a Zeiss Merlin SEM equipped with a X-Max 50 mm2 detector from Oxford Instruments. TEM imaging, EDX and SAED measurements were performed using a FEI Tanos F200S microscope, operated at 200 kV and equipped with a Thermo Fisher CETA camera. The VSM measurements were performed with a Lake Shore Cryotronics 7300 Series Vibrating Sample Magnetometer. Hysteresis loops were measured by applying the external field parallel to the surface of the thin film (i.e. within the easy plane set by shape anisotropy). Electron holography measurements were performed on a FEI Titan Themis microscope operated at 300 kV, equipped with an electron biprism and a Gatan K3 camera; -phase shifting off-axis holography method was used [66] and electrostatic contribution was removed to access the magnetic contribution.
@@
Supporting Information for XRD and XPS analysis, electron holography measurements, and micromagnetic simulations is available.
@@ Acknowledgements
The authors acknowledge the Center of Micro-NanoTechnology (CMi) and the Interdisciplinary Center for Electron Microscopy (CIME) of EPFL for their support during the development of the fabrication process and materials characterization, and the X-Ray Diffraction and Surface Analytics Facility for XPS measurements. A special gratitude goes to Dr. Rolf Allenspach (IBM Research-Zurich) for his valuable insights on the micromagnetic simulations, as well as to Dr. Floris Braakman, Liza Žaper, and prof. Martino Poggio from the University of Basel for their insightful discussions regarding the presented architecture. This work was supported as a part of NCCR SPIN, a National Centre of Competence in Research, funded by the Swiss National Science Foundation (grant number 225153).